Sally is 4 years older than Tim. Together the sum of their age is 20. How old is each?
Answer by EdenWolf(517)

Let x = Tim's age and x+4 = Sally's age. Set up an equation.
x+x+4=20
2x+4=20
2x=16
x=8
So Tim is 8 years old and Sally is 12 years old.